what is the magnitude of the electric field at a point of 30 cm to the right of a small sphere with a charge of -4.0 x 10-⁶C


1
Expert's answer
2021-05-18T11:10:33-0400

The magnitude of the electric field at this point is


"E=\frac{Q}{4\pi\epsilon_0r^2}=400\text{ kV/m}."